OpenStudy (anonymous):

the sum of the page numbers on the facing pagers of a book is 381. find the larger page number.

OpenStudy (anonymous):

x + x+ 1 = 381

OpenStudy (anonymous):

where x = smaller number

OpenStudy (anonymous):

the page numbers will be x and x+1, so x + (x+1) =381 2x+1=381 2x=380 x = 190 x+1 = 191 and this is the larger page number

OpenStudy (anonymous):

the page numbers will be x and x+1, so x + (x+1) =381 2x+1=381 2x=380 x = 190 x+1 = 191 and this is the larger page number

OpenStudy (anonymous):

beans1992 where did the 1 come in from?

OpenStudy (radar):

The facing pages will be consecutive, so one will be one more than the other.

OpenStudy (anonymous):

thank you radar!

OpenStudy (mertsj):